OpenStudy (anonymous):

What is the equation of a line passing through the point (a, b) and having a slope of b?

OpenStudy (anonymous):

The point-slope form of a line with a slope of m that passes through the point \[(x_o,y_o)\] is: \[y-y_o = m(x-x_o)\]

OpenStudy (anonymous):

I have y=-bx-b(1-a) y=bx-b(a-1) y=1/bx+b(a-1) y=-1/bx+b(1-a) :/ I dont know how to get it in the right form, it doesnt look like any point slope i have ever seen

OpenStudy (anonymous):

I cant see how they got any of those 4 options, i have no IDEA which one is correct

OpenStudy (anonymous):

If we plug in what we know, we'll get: \[y - b = b(x - a)\] \[y -b = bx-ab\] \[y=bx-ab+b = bx+b(1-a)\]

OpenStudy (anonymous):

Which is equivalent to: \[y=bx-b(a-1)\]

OpenStudy (anonymous):

oh! thank you! you rock!!! i have like 5 questions using this!